standard quality coffee sells for $18.00/kg and prime quality coffee sells for $24.00/kg. What quantity of each should be combined to produce 40 kg of blend to sell for $22.50/kg.
Answer by ewatrrr(24785)

x the amount of $24
24x + 18(40kg-x) = 22.5(40kg)
x = 4.5(40kg)/6, $24 variety. $18 variety 40kg-x